OpenStudy (anonymous):

Please help ASAP!! Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

hero (hero):

If 3, -13, and 5 + 4i are the zeroes of the polynomials, then x = 3 x = -13 x = 5 + 4i Note that complex zeroes come in conjugate pairs so there's another zero: x = 5 - 4i Move everything to the left side x - 3 = 0 x + 13 = 0 x - 5 - 4i = 0 x - 5 + 4i = 0 By zero product property: (x - 3)(x + 13)(x - 5 - 4i)(x - 5 + 4i) = 0

OpenStudy (anonymous):

@Hero then what is standard form?

hero (hero):

First multiply the factors together to find the polynomial expression.

OpenStudy (anonymous):

x^4-975

hero (hero):

Mind showing your work to demonstrate how you got that?

hero (hero):

I'm not sure I agree with your result

OpenStudy (anonymous):

I combined like terms

OpenStudy (anonymous):

@Hero